# Vision-Language Models for Medical Report Generation and Visual Question Answering: A Review

### v1

@@ 6 Medical Vision-Language Models
  • X-REM (Contrastive X-Ray REport Match) [Jeo+23] is a retrieval-based radiology RG model that uses an ITM score to measure the similarity of a chest X-ray image and radiology report for report retrieval. The VLM backbone of the model is ALBEF [Li+21]. ALBEF utilizes ViT-B/16 [Dos+21] as its image encoder and initializes the text encoder with the first 6 layers of the BERT [Dev+19] base model. The multimodal encoder in ALBEF, responsible for combining visual and textual features to generate ITM scores, is initialized using the final six layers of the BERT base model. X-REM leverages ALBEF’s pre-trained weights and performs further pre-training on X-rays paired with extracted impression sections ($2,192$ pairs), findings sections ($1,597$ pairs), or both ($2,192$ pairs) from the MIMIC-CXR [Joh+19a] dataset. Subsequently, the model is fine-tuned on the ITM task, where the scoring mechanism involves using the logit value for the positive class as the similarity score for image-text pairs. To address the positive skewness in medical datasets, 14 clinical labels obtained from the CheXbert [Smi+20] labeler are utilized. The model efficiently manages the computational burden associated with ITM scores by employing ALBEF’s pre-aligned unimodal embeddings. This involves narrowing down the candidate reports based on high cosine similarity with the input image before computing ITM scores. Additionally, the text encoder undergoes fine-tuning on natural language inference (NLI) task, utilizing datasets such as MedNLI [RS18] and RadNLI [Miu+21]. This step is crucial for preventing the retrieval of multiple reports with overlapping or conflicting information. X-REM achieves a BLUE-2 score of $0.186$ on the MIMIC-CXR (Findings only) dataset. The BERTScore of the model is $0.386$ on MIMIC-CXR (Findings only) and is $0.287$ on MIMIC-CXR (Impressions and Findings). The human evaluation of X-REM is described in Section 5.
@@
  • Visual Med-Alpaca [Shu+23] is a biomedical foundational model designed for addressing multimodal biomedical tasks like VQA. The model is constructed in the following way. First, image inputs go through a classifier to determine the appropriate module for transforming visual information into an intermediate text format. The currently supported modules include DePlot [Liu+22], utilized for interpreting plots and charts, and Med-GIT [Wan+22], fine-tuned specifically on the ROCO [Pel+18] dataset for understanding radiology images. The prompt manager then amalgamates textual information extracted from images and text inputs to construct the prompt for the LLM model, LLaMA-7B [Tou+23a]. However, before generating responses, LLaMa-7B undergoes both standard fine-tuning and LoRA [Hu+22] fine-tuning on a carefully curated set of $54,000$ medical question-answer pairs. The questions within this set are derived from question answering datasets such as MEDIQA QA [BSD19], MEDIQA RQE [BSD19], MedQA [Jin+21], MedDialog [Zen+20], and PubMedQA [Jin+19], with their corresponding answers synthesized using GPT-3.5-Turbo in the self-instruct [Wan+23a] manner. Human experts then meticulously filter and edit the obtained question-answer pairs to ensure quality and relevance. The evaluation of this model is still ongoing.
@@
  • CXR-RePaiR-Gen (Contrastive X-ray-Report Pair Retrieval based Generation) [Ran+23a] is a VLM designed for radiology RG that incorporates the RAG framework to mitigate the issue of hallucinated references in RG. The model leverages the pre-trained ALBEF [Lan+19] previously utilized in the CXR-ReDonE [RCR22] model. The ALBEF model consists of a ViT-B/16 [Dos+21] image encoder and the first 6 layers of BERT [Dev+19] as the text encoder, producing contrastively aligned image and text embeddings. Textual features are indexed in a vector database FAISS (Facebook AI Similarity Search), and when given a radiology image input, embeddings from the reports or sentences corpus with the highest dot-product similarity to the image embedding are retrieved. The CXR-PRO [RCR22] dataset is employed for text retrieval to gather relevant impressions for generating the radiology report. The retrieved impression sections from the CXR-PRO dataset serve as the context for the prompt to an LLM, along with instructions to generate the radiology report. Two distinct prompts are employed for generating free-text reports: one for the text-davinci-003 model and another for RG in a conversational setting with the GPT-3.5-Turbo and GPT-4 models. The model is evaluated on MS-CXR [Boe+22] and CXR-PRO datasets. There is no code provided for this model yet. CXR-RePaiR-Gen reaches a BERTScore score of $0.2865$ on the CXR-PRO dataset when based on GPT-4. Additionally, CXR-RePaiR-Gen achieves a score of $0.1970$ on MS-CXR when based on text-davinci-003. The model attains a RadGraph F1 score of $0.1061$ on the CXR-PRO dataset when based on GPT-4 and $0.0617$ on the MS-CXR dataset when it is based on text-davinci-003. In these instances, the CXR-RePaiR-Gen utilizes three retrieval samples per input during the RAG process.
@@
  • LLaVa-Med (Large Language and Vision Assistant for BioMedicine) [Li+23] is an adaptation of the VLM LLaVa [Liu+23a], specifically tailored for the medical domain through training on instruction-following datasets. The visual features are generated by the pre-trained CLIP ([Rad+21]) visual encoder ViT-L/14 [Dos+21]. The encoder can be substituted with BiomedCLIP [Zha+23a]. These features are passed through a linear projection layer, which converts them into tokens, and then, together with the tokenized instructions, are fed into the LLM LLaMa-7B [Tou+23a]. The LLM can be substituted with Vicuna [Chi+23]. After being initialized with the general-domain LLaVA, the model undergoes fine-tuning using curriculum learning. First, the model tries to understand and connect visual elements in biomedical images to the corresponding words or descriptions in the language model’s knowledge. To achieve that, a dataset consisting of $600,000$ image-caption pairs from the PMC-15 dataset, which was originally employed in the training of BiomedCLIP, is utilized. These image-caption pairs are transformed into an instruction-following dataset, where the instructions prompt the model to describe the corresponding image concisely or in detail. Given the language instruction and image input, the model is then prompted to predict the original caption. During this stage, the visual encoder and language model weights are kept frozen, with updates exclusively applied to the linear projection layer. The second stage of training focuses on aligning the model to follow diverse instructions. For this purpose, another instruction-following dataset is generated from PMC-15. For this dataset, instructions are designed to guide GPT-4 to generate multi-round questions and answers from the image caption and sentences from the original PMC paper that mentions the image [Li+23]. In this training phase, the model undergoes training on a set of $60,000$ images, each accompanied by its respective caption and multi-round questions and answers. Throughout this process, the weights of the visual encoder remain unchanged, preserving the previously acquired visual features. Meanwhile, the pre-trained weights of both the projection layer and the language model undergo continuous updates. This approach enables the model to effectively respond to a variety of instructions and perform well in generating dynamic and informative multi-round conversational content. Lastly, for VQA the model is fine-tuned and evaluated on VQA-RAD [Lau+18], SLAKE [Liu+21], and PathVQA [He+20] (see Table 3).
@@
  • CAT-ViL DeiT [BIR23] stands out as a specialized medical Vision-Language Model (VLM) tailored for Visual Question Answering (VQA) within surgical scenarios, with a unique focus on answer localization. The architecture incorporates a ResNet-18 [He+16] as the visual encoder, pre-trained on ImageNet [Den+09], and a customized pre-trained BERT tokenizer [See+22] for the text encoder. Central to its functionality is the Co-Attention gaTed Vision-Language (CAT-ViL) module, which enables interaction between visual and textual features and fuses them via a gating mechanism to obtain optimized multimodal embeddings. These features are then fused using a gating mechanism, yielding optimized multimodal embeddings. The model further integrates a pre-trained Data-efficient image Transformer (DeiT) [Tou+21] module to process these multimodal embeddings, aiming to acquire an optimal joint representation for comprehensive visual and textual understanding. In the context of VQA, the model adopts a standard classification head, while for answer localization within images, it employs the detection with transformers (DETR) [Car+20] head. The overall loss function comprises cross-entropy as the classification loss and L1-norm, along with the generalized intersection over union (GIoU) [Rez+19], serving as the localization loss. The model is trained on $1,560$ frames and $9,014$ QA pairs from the surgical datasets EndoVis 2018 [All+20]. The model achieved an accuracy of $61.92\%$ on the remaining data from EndoVis 2018 and $45.55\%$ on EndoVis 2017 [All+19] dataset.
@@
  • MUMC (Masked image and text modeling with Unimodal and Multimodal Contrastive losses) [Li+23a] is a VLM designed for medical VQA. The model utilizes a ViT-B/12 [Dos+21] as its image encoder, the first 6 layers of BERT [Dev+19] as its text encoder, and the last 6 layers of BERT as its multimodal encoder. The multimodal encoder incorporates cross-attention layers to align visual and textual features. For pre-training, the model employs a combination of contrastive learning, MLM, and ITM objectives. Also, the model utilizes a newly introduced masked image strategy, randomly masking 25% of image patches as a data augmentation technique. This exposes the model to a greater variety of visual contexts and enables learning representations that are more robust to partially occluded inputs. The pre-training is performed on the ROCO [Rad+21], MedICaT [Sub+20], and ImageCLEF (Image Retrieval in Cross-Language Evaluation Forum) Caption [Rüc+22] datasets. For downstream VQA tasks, an answering decoder is added on top of the multimodal encoder to generate answer text tokens. The encoder weights are initialized from pre-training, and the model is fine-tuned and evaluated on VQA-RAD [Lau+18], SLAKE [Liu+21], and PathVQA [He+20] (see Table 3).
@@
  • Med-Flamingo [Moo+23] is a multimodal few-shot learner model based on the Flamingo [Ala+22] architecture, adapted to medical domain. The model is pre-trained on the MTB [Moo+23] dataset, a newly curated collection comprising $4,721$ segments from various Medical TextBooks, encompassing both textual content and images. Each segment is designed to contain at least one image and up to $10$ images, with a specified maximum length. Also, it is pre-trained on $1.3$ M image-caption pairs from the PMC-OA [Lin+23] dataset. The Model’s few-shot capabilities are achieved through training on these mixed text and image datasets, enabling it to generalize and perform diverse multimodal tasks with only a few examples. The model utilizes pre-trained frozen CLIP vision encoder ViT-L/14 for visual features generation. To convert these visual features into a fixed number of tokens, the model employs a module known as the perceiver resampler, which is trained from scratch. Subsequently, these tokens, along with tokenized text inputs, undergo further processing in a pre-trained frozen LLM LLaMA-7B [Tou+23a], enhanced with strategically inserted gated cross-attention layers that are also trained from scratch. This augmentation not only facilitates the learning of novel relationships but also bolsters training stability. The model’s performance is evaluated on established benchmarks such as VQA-RAD [Lau+18] and PathVQA [He+20], demonstrating its effectiveness in medical visual question-answering. The exact match scores for MedFlamingo demonstrate a few-shot performance of $0.200$ on VQA-RAD and $0.303$ on PathVQA. In contrast, the zero-shot performance yields an exact match score of $0.000$ on VQA-RAD and $0.120$ on PathVQA. Additionally, it is evaluated on a specifically created Visual USMLE (United States Medical Licensing Examination) dataset, comprising $618$ challenging open-ended USMLE-style questions augmented with images, case vignettes, and tables of laboratory measurements, covering a diverse range of medical specialties. The human evaluation of the MedFlamingo model on VQA-RAD, PathVQA, and Visual USMLE datasets is described in Section 5.
@@
  • RaDialog [Pel+23] is a VLM that integrates automated radiology RG with conversational assistance. The model incorporates BioViL-T [Ban+23], a hybrid model that amalgamates the strengths of ResNet-50 [He+16] and Transformer [Vas+17] architectures. Pre-trained on radiology images and reports, BioViL-T serves as a vision encoder that generates patch-wise visual features. The extracted features undergo alignment through a BERT [Dev+19] model, transforming them into a concise representation of $32$ tokens. The model incorporates the CheXpert classifier to offer organized findings in medical images. These findings are generated based on labels obtained from the CheXbert [Smi+20] model. The classifier is trained independently using labels predicted by CheXbert from the findings section of radiology reports. The model integrates visual features, structured findings, and the directive “Write a radiology report” into a singular prompt, which is used as input for the LLM, a Vicuna-7B [Chi+23] model fine-tuned using LoRA [Hu+22]. The training is performed on X-ray image-report pairs from MIMIC-CXR [Joh+19a] dataset. RaDialog achieves a BLUE-4 score of $0.095$, ROUGE-L score of $0.2710$, METEOR score of $0.14$, and BERTScore of $0.400$ on the MIMIC-CXR dataset. To address the challenge of catastrophic forgetting during training and ensure the model’s capability across diverse downstream tasks, it is specifically trained on the newly created Instruct [Pel+23] dataset. This dataset is meticulously curated to encompass a spectrum of 8 diverse tasks: RG, NLE, complete CheXpert QA, binary CheXpert QA, region QA, summarization, report correction, and reformulation report using simple language. Each task is accompanied by carefully formulated prompts, tailored to elicit specific responses from the model. For instance, some prompts involve answering questions about particular X-ray regions. RaDialog trained on the Instruct dataset achieves an F1 score of 0.397 on the binary CheXpert QA task and 0.403 on the complete CheXpert QA task while RaDialog without being trained on Instruct achieves lower F1 scores of 0.018 and 0.098, respectively.
Model SLAKE VQA-RAD PathVQA VQA-Med VQA-Med
open close open close open close 2019 2021
-ended -ended -ended -ended -ended -ended
@@
[Li+23a]
7 Challenges & Future Directions

### v2

@@ 4.3.7 Contrastive X-Ray REport Match (X-REM)
X-REM is a retrieval-based radiology RG model that uses an ITM score to measure the similarity of a chest X-ray image and radiology report for report retrieval [jeong2023xrem]. The VLM backbone of the model is ALBEF [li2021albef]. ALBEF utilizes ViT-B/16 [dosovitskiy2021vit] as its image encoder and initializes the text encoder with the first 6 layers of the BERT [devlin2019bert] base model. The multimodal encoder in ALBEF, responsible for combining visual and textual features to generate ITM scores, is initialized using the final six layers of the BERT base model. X-REM leverages ALBEF’s pre-trained weights and performs further pre-training on X-rays paired with extracted impression sections ($2,192$ pairs), findings sections ($1,597$ pairs), or both ($2,192$ pairs) from the MIMIC-CXR [johnson2019mimiccxr] dataset. Subsequently, the model is fine-tuned on the ITM task, where the scoring mechanism involves using the logit value for the positive class as the similarity score for image-text pairs. To address the positive skewness in medical datasets, 14 clinical labels obtained from the CheXbert [smit2020chexbert] labeler are utilized. The model efficiently manages the computational burden associated with ITM scores by employing ALBEF’s pre-aligned unimodal embeddings. This involves narrowing down the candidate reports based on high cosine similarity with the input image before computing ITM scores. Additionally, the text encoder undergoes fine-tuning on natural language inference (NLI) task, utilizing datasets such as MedNLI [romanov2018mednli] and RadNLI [miura2021radnli]. This step is crucial for preventing the retrieval of multiple reports with overlapping or conflicting information. X-REM achieves a BLEU-2 score of $0.186$ on the MIMIC-CXR (Findings only) dataset. The BERTScore of the model is $0.386$ on MIMIC-CXR (Findings only) and is $0.287$ on MIMIC-CXR (Impressions and Findings). The human evaluation of X-REM is described in Section 4.2.
4.3.8 Visual Med-Alpaca
@@
Visual Med-Alpaca is a biomedical foundational model designed for addressing multimodal biomedical tasks like VQA [shu2023vismedalpaca]. The model is constructed in the following way. First, image inputs go through a classifier to determine the appropriate module for transforming visual information into an intermediate text format. The currently supported modules include DePlot [liu2023deplot], utilized for interpreting plots and charts, and Med-GIT [wang2022git], fine-tuned specifically on the ROCO [pelka2018roco] dataset for understanding radiology images. The prompt manager then amalgamates textual information extracted from images and text inputs to construct the prompt for the LLM model, LLaMA-7B [touvron2023llama]. However, before generating responses, LLaMa-7B undergoes both standard fine-tuning and LoRA [hu2022lora] fine-tuning on a carefully curated set of $54,000$ medical question-answer pairs. The questions within this set are derived from question-answering datasets such as MEDIQA QA [benabacha2019mediqa], MEDIQA RQE [benabacha2019mediqa], MedQA [jin2021medqa], MedDialog [zeng2020meddialog], and PubMedQA [jin2019pubmedqa], with their corresponding answers synthesized using GPT-3.5-Turbo in the self-instruct [wang2023selfinstruct] manner. Human experts then meticulously filter and edit the obtained question-answer pairs to ensure quality and relevance. The evaluation of this model is still ongoing [shu2023vismedalpaca].
4.3.9 Contrastive X-ray-Report Pair Retrieval based Generation (CXR-RePaiR-Gen)
@@
CXR-RePaiR-Gen is designed for radiology RG that incorporates the RAG framework to mitigate the issue of hallucinated references [ranjit2023cxrrepairgen]. The model leverages the pre-trained ALBEF [lan2019albert] previously utilized in CXR-ReDonE [ramesh2022cxrredone]. The ALBEF model consists of a ViT-B/16 [dosovitskiy2021vit] image encoder and the first 6 layers of BERT [devlin2019bert] as the text encoder, producing contrastively aligned image and text embeddings. Textual features are indexed in a vector database, Facebook AI Similarity Search (FAISS). When given a radiology image input, embeddings from the reports or sentences corpus with the highest dot-product similarity to the image embedding are retrieved. The CXR-PRO [ramesh2022cxrredone] dataset is employed for text retrieval to gather relevant impressions for generating the radiology report. The retrieved impression sections from the CXR-PRO dataset serve as the context for the prompt to an LLM, along with instructions to generate the radiology report. Two distinct prompts are employed for generating free-text reports: one for the text-davinci-003 model and another for RG in a conversational setting with the GPT-3.5-Turbo and GPT-4 models. The model is evaluated on MS-CXR [boecking2022mscxr] and CXR-PRO datasets. There is no code provided for this model yet. CXR-RePaiR-Gen reaches a BERTScore score of $0.2865$ on the CXR-PRO dataset when based on GPT-4. Additionally, CXR-RePaiR-Gen achieves a score of $0.1970$ on MS-CXR when based on text-davinci-003. The model attains a RadGraph F1 score of $0.1061$ on the CXR-PRO dataset when based on GPT-4 and $0.0617$ on the MS-CXR dataset when it is based on text-davinci-003. In these instances, the CXR-RePaiR-Gen utilizes three retrieval samples per input during the RAG process.
4.3.10 Large Language and Vision Assistant for BioMedicine (LLaVa-Med)
@@
LLaVa-Med is an adaptation of the VLM LLaVa [liu2023llava], specifically tailored for the medical domain through training on instruction-following datasets [li2023llavamed]. The visual features are generated by the pre-trainedCLIP [radford2021clip] visual encoder ViT-L/14 [dosovitskiy2021vit]. The encoder can be substituted with BiomedCLIP [zhang2023biomedclip]. These features are passed through a linear projection layer, which converts them into tokens, and then, together with the tokenized instructions, are fed into the LLM LLaMa-7B [touvron2023llama]. The LLM can be substituted with Vicuna [chiang2023vicuna]. After initializing with the general-domain LLaVA, the model undergoes fine-tuning using curriculum learning. First, the model tries to understand and connect visual elements in biomedical images to the corresponding words or descriptions in the language model’s knowledge. To achieve that, a dataset consisting of $600,000$ image-caption pairs from the PMC-15 dataset, which was originally employed in the training of BiomedCLIP, is utilized. These image-caption pairs are transformed into an instruction-following dataset, where the instructions prompt the model to describe the corresponding image concisely or in detail. Given the language instruction and image input, the model is then prompted to predict the original caption. During this stage, the visual encoder and language model weights are kept frozen, with updates exclusively applied to the linear projection layer. The second stage of training focuses on aligning the model to follow diverse instructions. For this purpose, another instruction-following dataset is generated from PMC-15. For this dataset, instructions are designed to guide the GPT-4 model to generate multi-round questions and answers from the image caption and sentences from the original PMC paper that mentions the image [li2023llavamed]. In this training phase, the model undergoes training on a set of $60,000$ images, each accompanied by its respective caption and multi-round questions and answers. Throughout this process, the weights of the visual encoder remain unchanged, preserving the previously acquired visual features. Meanwhile, the pre-trained weights of both the projection layer and the language model undergo continuous updates. This approach enables the model to effectively respond to a variety of instructions and perform well in generating dynamic and informative multi-round conversational content. Lastly, for VQA, the model is fine-tuned and evaluated on VQA-RAD [lau2018vqarad], SLAKE [liu2021slake], and PathVQA [he2020pathvqa] (see Table 3).
4.3.11 XrayGPT
@@
4.3.12 Co-Attention gaTed Vision-Language Data-efficient image Transformer (CAT-ViL DeiT)
@@
CAT-ViL DeiT stands out as a specialized VLM tailored for VQA within surgical scenarios, with a unique focus on answer localization [bai2023catvil]. The architecture incorporates a ResNet-18 [he2016resnet] as the visual encoder, pre-trained on ImageNet [deng2009imagenet], and a customized pre-trained BERT tokenizer [seenivasan2022surgicalvqa] for the text encoder. Central to its functionality is the Co-Attention gaTed Vision-Language (CAT-ViL) module, which enables interaction between visual and textual features and fuses them via a gating mechanism to obtain optimized multimodal embeddings. These features are then fused using a gating mechanism, yielding optimized multimodal embeddings. The model further integrates a pre-trained Data-efficient image Transformer (DeiT) [touvron2021deit] module to process these multimodal embeddings, aiming to acquire an optimal joint representation for comprehensive visual and textual understanding. In the context of VQA, the model adopts a standard classification head, while for answer localization within images, it employs the detection with transformers (DETR) [carion2020detr] head. The overall loss function comprises cross-entropy as the classification loss and L1-norm, along with the generalized intersection over union (GIoU) [rezatofighi2019giou], serving as the localization loss. The model is trained on $1,560$ frames and $9,014$ QA pairs from the surgical datasets EndoVis 2018 [allan2020endovis18]. The model achieved an accuracy of $61.92\%$ on the remaining data from EndoVis 2018 and $45.55\%$ on EndoVis 2017 [allan2019endovis17] dataset.
4.3.13 Masked image and text modeling with Unimodal and Multimodal Contrastive losses (MUMC)
@@
MUMC utilizes a ViT-B/12 [dosovitskiy2021vit] as its image encoder, the first 6 layers of BERT [devlin2019bert] as its text encoder, and the last 6 layers of BERT as its multimodal encoder [li2023mumc]. The multimodal encoder incorporates cross-attention layers to align visual and textual features. For pre-training, the model employs a combination of contrastive learning, MLM, and ITM objectives. Also, the model utilizes a newly introduced masked image strategy, randomly masking 25% of image patches as a data augmentation technique. This exposes the model to a greater variety of visual contexts and enables learning representations that are more robust to partially occluded inputs. The pre-training is performed on the ROCO [radford2021clip], MedICaT [subramanian2020medicat], and Image Retrieval in Cross-Language Evaluation Forum (ImageCLEF) caption [ruckert2022imageclef] datasets. For downstream VQA tasks, an answering decoder is added on top of the multimodal encoder to generate answer text tokens. The encoder weights are initialized from pre-training, and the model is fine-tuned and evaluated on VQA-RAD [lau2018vqarad], SLAKE [liu2021slake], and PathVQA [he2020pathvqa] (see Table 3).
4.3.14 Med-Flamingo
@@
Med-Flamingo is a multimodal few-shot learner model based on the Flamingo [alayrac2022flamingo] architecture, adapted to the medical domain [moor2023medflamingo]. The model is pre-trained on the MTB [moor2023medflamingo] dataset, a newly curated collection comprising $4,721$ segments from various Medical TextBooks, encompassing both textual content and images. Each segment is designed to contain at least one image and up to $10$ images, with a specified maximum length. Also, it is pre-trained on $1.3$ M image-caption pairs from the PMC-OA [lin2023pmcclip] dataset. The Model’s few-shot capabilities are achieved through training on these mixed text and image datasets, enabling it to generalize and perform diverse multimodal tasks with only a few examples. The model utilizes a pre-trained frozen CLIP vision encoder ViT-L/14 for visual feature generation. To convert these visual features into a fixed number of tokens, the model employs a module known as the perceiver resampler, which is trained from scratch. Subsequently, these tokens, along with tokenized text inputs, undergo further processing in a pre-trained frozen LLM LLaMA-7B [touvron2023llama], enhanced with strategically inserted gated cross-attention layers that are also trained from scratch. This augmentation not only facilitates the learning of novel relationships but also bolsters training stability. The model’s performance is evaluated on established benchmarks such as VQA-RAD [lau2018vqarad] and PathVQA [he2020pathvqa], demonstrating its effectiveness in medical visual question-answering. The exact match scores for MedFlamingo demonstrate a few-shot performance of $0.200$ on VQA-RAD and $0.303$ on PathVQA. In contrast, the zero-shot performance yields an exact match score of $0.000$ on VQA-RAD and $0.120$ on PathVQA. Additionally, it is evaluated on a specifically created Visual United States Medical Licensing Examination (USMLE) dataset, comprising $618$ challenging open-ended USMLE-style questions augmented with images, case vignettes, and tables of laboratory measurements, covering a diverse range of medical specialties. The human evaluation of the Med-Flamingo model on VQA-RAD, PathVQA, and Visual USMLE datasets is described in Section 4.2.
4.3.15 RaDialog
@@
RaDialog is a VLM that integrates automated radiology RG with conversational assistance [pellegrini2023radialog]. The model incorporates BioViL-T [bannur2023biovilt], a hybrid model that fuses the strengths of ResNet-50 [he2016resnet] and Transformer [vaswani2017attention] architectures. Pre-trained on radiology images and reports, BioViL-T serves as a vision encoder that generates patch-wise visual features. The extracted features undergo alignment through a BERT [devlin2019bert] model, transforming them into a concise representation of $32$ tokens. The model incorporates the CheXpert classifier to offer organized findings in medical images. These findings are generated based on labels obtained from the CheXbert [smit2020chexbert] model. The classifier is trained independently using labels predicted by CheXbert from the findings section of radiology reports. The model integrates visual features, structured findings, and the directive “Write a radiology report” into a singular prompt, which is used as input for the LLM, a Vicuna-7B [chiang2023vicuna] model fine-tuned using LoRA [hu2022lora]. The training is performed on X-ray image-report pairs from MIMIC-CXR [johnson2019mimiccxr] dataset. RaDialog achieves a BLEU-4 score of $0.095$, ROUGE-L score of $0.2710$, METEOR score of $0.14$, and BERTScore of $0.400$ on the MIMIC-CXR dataset. To address the challenge of catastrophic forgetting during training and ensure the model’s capability across diverse downstream tasks, it is specifically trained on the newly created Instruct [pellegrini2023radialog] dataset. This dataset is meticulously curated to encompass a spectrum of 8 diverse tasks: RG, NLE, complete CheXpert QA, binary CheXpert QA, region QA, summarization, report correction, and reformulation report using simple language. Carefully formulated prompts accompany each task, tailored to elicit specific responses from the model. For instance, some prompts involve answering questions about particular X-ray regions. RaDialog trained on the Instruct dataset achieves an F1 score of $0.397$ on the binary CheXpert QA task and $0.403$ on the complete CheXpert QA task. In contrast, RaDialog without being trained on Instruct achieves lower F1 scores of $0.018$ and $0.098$, respectively.
Model SLAKE SLAKE VQA-RAD VQA-RAD PathVQA PathVQA
open close open close open close VQA-Med VQA-Med
-ended -ended -ended -ended -ended -ended 2019 2021
@@
[li2023mumc]
5 Challenges and Potential Future Directions